OpenStudy (anonymous):

Given the function g(x) = −4x + 5, find g(−4).

OpenStudy (anonymous):

g(−4).=-4(-4)+5=21

OpenStudy (anonymous):

how bout this one Given the function f(x) = 2(x + 6), find x if f(x) = 22.

OpenStudy (anonymous):

Put x = -4 in the the function: g(-4) = -4(-4) + 5 g(-4) = 21...

OpenStudy (anonymous):

Then put f(x) = 22 S0, 22 = 2(x + 6) 11 = x + 6 x = 5 is the answer...

OpenStudy (anonymous):

i dont get wat happened to the 2

OpenStudy (anonymous):

In simple way, divide by 2 both the sides, 22/11 = 2(x + 6)/2 So, the 2 will get cancelled leaving 11 = x + 6.. Getting or not??

OpenStudy (anonymous):

ok cool